# Measuring Value Expressions in Social Media Posts

### v2

@@ Abstract
The value alignment of sociotechnical systems has become a central debate but progress in this direction requires the measurement of the expressions of values. While the rise of large-language models offer new possible opportunities for measuring expressions of human values (e.g., humility or equality) in social media data, there remain both conceptual and practical challenges in operationalizing value expression in social media posts: what value system and operationalization is most applicable, and how do we actually measure them? In this paper, we draw on the Schwartz value system as a broadly encompassing and theoretically grounded set of basic human values, and introduce a framework for measuring Schwartz value expressions in social media posts at scale. We collect 32,370 ground truth value expression annotations from N=1,079 people on 5,211 social media posts representative of real users’ feeds. We observe low levels of inter-rater agreement between people, and low agreement between human raters and LLM-based methods. Drawing on theories of interpretivism — that different people will have different subjective experiences of the same situation — we argue that value expression is (partially) in the eye of the beholder. In response, we construct a personalization architecture for classifying value expressions. We find that a system that explicitly models these differences yields predicted value expressions that people agree with more than they agree with other people. These results contribute new methods and understanding for the measurement of human values in social media data.
@@ Introduction
Social media provides a rich and complex substrate to understand patterns of human life, behavior and discourse (lazer2009computational). The rise of large language models (LLMs) offers new opportunities for measuring a wide variety of constructs in social media data (ziems2024can; jia2024embedding). One construct of interest for social media, growing out of the focus on value alignment of sociotechnical systems (russell2015research), is human values: the core beliefs that serve as guiding principles and in turn shape decisions, behavior, and perception (kolluri2025alexandria; bernstein2023embedding; stray2024building). Values expressed on social media, therefore, offer a reflection of discourse and sentiment, that in turn affect discourse and behavior. The possibility of measuring human values in social media posts opens new opportunities to build a better understanding of how discourse operates online, and to design the next generation of social media algorithms and platforms. Past work has shown how values and related moral priorities explain patterns of belief and behavior in domains as varied as vaccine hesitancy (weinzierl2022hesitancy), pro-environmental behavior (schultz1998values), social norms (forbes2020social), prosociality (heilman2020personal), human development (inglehart2020modernization), consumer behavior (krystallis2012usefulness; stathopoulou2019effect), and news story framing (mokhberian2020moral). Beyond understanding, the capability to measure values could also enable algorithmic objectives for feed-based ranking that support both individual and societal goals. In particular, there is growing interest in aligning ranking algorithms with societal values (bail2022breaking; bernstein2023embedding; stray2024building; kolluri2025alexandria), focusing on values such as democratic process (jia2024embedding; piccardi2024social), well-being (stray2020aligning), and downranking low-quality information (epstein2020will).
@@
Before these opportunities can be realized, there remain both conceptual and practical challenges in operationalizing and measuring value expressions in social media. By value expressions, we mean the extent to which a social media post expresses attributes associated with a given value. Identifying value expression can be straightforward when a post directly references the value: for example, the value of equality is clearly expressed in the post “Not all disabilities are visible!” (Figure 1). Likewise, that same post clearly does not contain an expression of the value of tradition.
@@
Figure 1: An example social media post for value annotation. All annotators in our dataset agree that this post strongly expresses Universal Concern, and they all agree that the post does not express Tradition—but, annotators disagree substantially on whether the post expresses Humility. In this paper, we model these interpersonal differences for more accurate value classification.
@@
However, not in all cases are value expressions are as straightforward to classify on social media. The core issue is that value expression is a fundamentally subjective construct, leading to different labels from different people (van2023differences; basile2021we; orlikowski2023ecological). For example, while the Figure 1 post that “Not all disabilities are visible!” clearly is an expression of equality and clearly is not an expression of tradition, is it an expression of the value of humility? For some, the post strongly expresses a reminder to be humble about the conclusions we draw about other people’s (lack of) visible disabilities. For others, their perception of that value expression is much weaker. Both interpretations are valid, and shaped by an individual’s background and how they both conceptualize these values and understand the social media post itself. As a result, we argue that there exists conceptual and methodological ambiguity about how value expression is operationalized. While existing work (e.g., qiu2022valuenet; van2023differences; borenstein2024investigating) has developed classifiers for detecting basic human values, these models have presumed there exists in all cases one unambiguous correct answer to be learned.
@@
In this paper, we develop methods of identifying value expression in social media posts under an assumption of subjectivity in labeling value expressions. In particular, we collect a large annotated dataset where 1,079 people redundantly annotate values in 5,211 posts. To do so, we use a well-studied and theoretically grounded value system drawn from cross-cultural psychology - the Schwartz value system (schwartz2022measuring; schwartz1992universals) - that offers a grounded, universal set of basic human values (although we note that our approach can be adapted to other value systems as well). We evaluate to what extent large language models (LLMs) can identify the expression of values in these posts, comparing off-the-shelf LLM models with fine-tuned and personalized models, and both to individuals and groups of annotators.
@@
On this task, we find that not only are there low levels of agreement between people, but also that an off-the-shelf LLM exhibits ever lower levels of agreement with people than people’s levels of agreement with each other. To make sense of these findings, we draw on the theory of interpretetivism — that different people will have different subjective experiences of the same situation (soden2024evaluating; berger2016social) — to suggest that people’s divergent perceptions of both the posts and values is driving the low levels of human-human agreement we observe, as well as the lack of performance of the base GPT model.
@@
With this idea in mind, we implement a system that explicitly models interpersonal differences by fine-tuning a large model and incorporating the recommendation system inspired techniques of utilizing a cold start dataset for each user to make predictions. We find that our personalized approach results in a performant model that individuals agree with more than the rate at which they agree with other people (66% relative improvement in $\rho$), or with the consensus vote (28% relative improvement in $\rho$).
Our corresponding contributions are:
  1. 1.
    Demonstration that value expression exhibits high levels of disagreement, which underscores the perils of relying on majority aggregated labels in this task,
  2. 2.
    A theoretically-grounded personalization framework for calibrating predictions from a fine-tuned large language model to an individual based on their own responses that people agree with more than other people or LLMs with few-shot prompting, and
  3. 3.
    A dataset of fine-grained value expression annotations available to researchers for future work.
@@
Figure 2: The refined Schwartz value system. The system is hierarchically nested into a highest level: Outcomes for Self (right) vs Outcomes for Others (left), a high level: Self-Transcendence (top left), Conservation (bottom left), Self-Enhancement (bottom right) and Openness to Change (top left), and 19 low-level values. Adapted from (schwartz2012refining)
@@
To operationalize values, we employ Schwartz’s theory of Basic Human Values (schwartz2012refining), a theory from Cultural Psychology for mapping individual values. Schwartz defined (basic) values as “trans-situational goals, varying in importance, that serve as guiding principles in the life of a person or group” and sought to identify a set of values that could be recognized in all societies. The original theory (schwartz1992universals) identified 10 basic values from a 56-item questionnaire administered to 9140 students and teachers across 20 countries. The refined theory identified 19 values that exist on a circular continuum with 57-item questionnaire administered to 6,059 participants across 10 countries (see Figure 2). A core assumption of Schwartz’s theory of Basic Human Values is that: “The circular arrangement of values represents a continuum of related motivations, like the circular continuum of colors, rather than a set of discrete motivations” (davidov2008bringing). This circumplex structure is also reflected in the hierarchical branching structure of the values, which starts at the highest level (Outcomes for Self vs Others), then branches further to four high-level values (Openness to Change, Self-Enhancement, Self-Transcendence, and Conservation), and then 19 low-level values. Values close to each other on the wheel are theorized to have similar and compatible underlying motivations, while values on opposite sides are theorized to be in tension. While there are reasons to critique the Schwartz Value System (see e.g. chakroff2015discovering), the theory ultimately offers conceptual precision, depth, and breadth, as well as offering more theoretical plausibility, analytical utility, and empirical grounding than political values (goren2020values).
Automatically Detecting Schwartz Values
@@
qiu2022valuenet create a dataset of human attitudes on over 20,000 text scenarios called ValueNet and a transformer-based regression model for dialogue tasks. kiesel2022identifying create a dataset of the values behind arguments from four geographical cultures. ponizovskiy2020development developed a dictionary of words for Schwartz values based on a corpus of Facebook updates, blog posts, essays, and book chapters. boyd2015values use free-response survey methods to compare individual values reconstructed from text with traditional survey instruments, showing advancements in understanding in both the structure and content of values van2023differences use data from qiu2022valuenet and kiesel2022identifying to train a model for predicting value expression in Reddit posts, which they apply to a dataset of 11.4M comments from 19K users to construct value profiles of individual users. They find that dissimilarity of users’ values is associated with the levels of disagreement between those users.
borenstein2024investigating train value extraction models to classify both the presence and polarity of Schwartz values in reddit posts. They identify patterns like a strong negative stance towards conformity in subreddits such as r/Vegan and r/AbolishTheMonarchyand an overall correlation between tradition values and the conservativeness of U.S. states.
@@
This body of work uniformly makes the assumption that there exists an unambiguous “ground truth” value label, neglecting the inherent subjectivity of the task. In the next subsection, we discuss work that explores an alternative, interpretivist stance.
Interpretivism and Personalization
@@
The work highlighted above on automatically detecting Schwartz values offers an exciting lens on the values represented in text. However, in every case, these models assume a fixed meaning for both the content and the values, and a corresponding unambiguous “objective” value label in each case. This is in line with the “longstanding” paradigm in data labeling (fleisig2024perspectivist) which seeks to collect “ground truth” labels (nowak2010reliable; snow2008cheap) by aggregating labels from annotators. An alternative, the perspectivist paradigm (basile2021we; fleisig2024perspectivist; plank2022problem; mcdonald2019reliability; frenda2025perspectivist) argues that variation in annotator labels is not purely a source of error to be expunged, but rather is intrinsically meaningful. We argue that in our task of identifying the expression of basic human values in social media posts, there is no singularly correct “ground truth” label that everyone agrees with.
An alternative but complementary lens is personalization, which seeks to build models responsive to heterogeneous preferences. Past work in the context of aligning LLMs has explored personalized reward models (chen2024pad; jang2023personalized; li2024personalized), and aligning to users’ beliefs and opinions (hwang2023aligning; do2025aligning; suh2025language).
Value Systems and Their Measurement
There has been increased discussion and research on measuring values in textual data. Across the AI alignment literature, there exist hundreds of potential values (sorensen2024value; huang2025values) that AI systems could conform to, offering a large platter of categorical options but few approaches that relate these values to each other in a systematic way. In the context of social media, nguyen2024measuring leverage moral foundations theory (MFT) (haidt2012righteous) to fine-tune a large language model to these concepts in news media and online discussions. They show that these moral foundations can be used to meaningfully describe people’s stances on a variety of social issues, such as climate change, abortion, and religion.
@@
In this section, we first report our methods for curating a dataset of social media posts. Next, we discuss our methods for annotating this dataset for values. Finally, in the next section, we report our approach to modeling and evaluating human and AI-based value expressions.
@@
We recruited a sample of social media users (N=281) quota-matched to be a representative sample of the US population on age, ethnicity, gender and partisanship using Prolific in May 2024. Participants were all from the United States and were required to be 18 years old or older. We were unable to meet the quota for participants in the age range 55-100 (92/109) so we upsample posts from participants in this age range by 18.4% when constructing our silverset of social media posts. Within this sample, the median age was 43 (min=19, max=80), 58.36% White/Caucasian, 51.2% female and 30.6% Democrat, 28.8% Republican and 40.6% Independent. Participants were paid $\sim\mathdollar 15.42$ an hour and took approximately 7 minutes to complete the study.
@@
After they consented to the research, participants were asked to install a Chrome Extension that downloaded posts from their X (Twitter) feeds, both their algorithmically curated For You page (FYP) and their in-network Following feeds simultaneously. We collect and annotate posts from both the Following feed and FYP to account for any systematic differences that may exist between them, and to pave the way for work that uses both to measure algorithmic amplification (see Discussion).
When a browser opens the X page, tweets are fed to the browser client in “batches”: a single API request returns a set of tweets at a time. The Chrome Extension aims to download 16 batches of tweets: 7 from the FYP, and 9 from the Following page. Each batch for FYP consists of approximately 30 tweets, whereas the Following feed returns a variable number of tweets up to 100. From the 281 participants, we collected 265,442 posts, with an average of 207 posts from the FYP and 776 from the Following page per user. Among these 265,442 posts, there are 151,740 posts with unique, non-empty text in English (using Google’s Chromium language detection tool cld2). We also screen out private posts (n=6,025; 3.97%), deleted posts (n=2,004; 1.32%) and/or reply tweets that are replying to posts not displayed in the user’s feed (n=7,581; 4.99%) to focus only on the 142,652 public posts we can render with proper context.
@@
We next prepare these posts for human annotation, since some posts are either inappropriate (NSFW) to show to annotators or incomprehensible to most annotators. We filter the posts for comprehensibility (grounded in the constructs of readability, coherence, spam behavior, and context required for understanding) and NSFW (to comply with our institution’s IRB for data annotation) using an automated GPT-based pipeline, with the prompt that rates all 19 values jointly each on a 0-6 scale based on a short description (see the Appendix for full text) with GPT-4o, temperature=1.0 and seed=0. We filter out any posts that score less than 3 on a comprehensibility scale, resulting in a dataset of 89,383 (62.6%) posts. We then filter out 1,937 posts with NSFW content, resulting in a final dataset of 87,446 posts.
@@
The occurrence rate of different values in social media posts varies, as do the numbers of posts per person across both Following and FYP feeds. Therefore, uniform sampling procedure may yield a sample that underrepresents rarer values and overrepresents certain users for annotation. To construct a final set of posts for human annotation, we stratify on user and post source as well as on a preliminary LLM-based measure of value expression tuned for recall over precision to prioritize catching as many potential value-expressing posts as possible. Since stratification requires an initial estimate of which value a post may express, we first built an automated classifier tuned for recall rather than precision—that is, we prioritized sampling as many potential value-expressing posts as possible, even if this meant including some false positives.
@@
In particular, we perform a preliminary rough classification of these posts for all 19 of the Basic Human Values (schwartz2012refining) on a 7-point Likert scale [0-6] using another automated GPT-based pipeline, with the prompt defined in Table A4 with GPT-4o, temperature=1.0 and seed=0. This allows us to upsample posts that may reflect rarer values. For this upsampling, we say a given post reflects a given value if the GPT annotation pipeline scored it a 4 or above. We then sample a final set of 5,227 posts for human annotation by stratifying across users, post source (FYP vs following), and value expression. In particular, for each user and each value, we sample a tweet from that user’s feed that reflects that value. We sample FYP and Following posts with equal probability, but for a given user, if only one type of post exists, we return that tweet. Because we did not meet our quota for participants over the age of 55 (by 17 participants), we randomly sample (without replacement) 17 of the 76 participants over the age of 55 who completed the survey and repeat this process for each of them to ensure representation of posts from users over the age of 55. Of the 5,227, 3,182 came from users’ Following feeds (60.8%) and 2,045 (39.1%) came from users’ FYP.
Annotating Social Media Posts for Basic Human Values
@@
This process yielded a dataset of 5,227 posts that we can use to study human value annotation and to predict the values expressed in those posts. In this section, we describe our ground truth data annotation method.

### v3

# Whose Values? Measuring the (Subjective) Expression of Basic Human Values in Social Media Posts

@@ Abstract
The value alignment of sociotechnical systems has become a central debate, but progress depends on how human values are perceived in the content these systems surface and how such perceptions can be measured at scale. Social media platforms are a prominent class of sociotechnical systems where algorithmic curation shapes exposure to value-laden content at scale. Large-language models offer new opportunities for measuring expressions of human values (e.g., humility or equality) in social media data, but value expressions can be subjective: different people will annotate the same post with different values. In this paper, we draw on the Schwartz value system as a broadly encompassing and theoretically grounded set of basic human values, and introduce a framework to personalize the measurement of expressions of Schwartz values in social media posts at scale. We collect 32,370 ground truth value expression annotations from N=1,079 people on 5,211 social media posts representative of real users’ feeds. Due to the subjectivity of the task, we observe low levels of inter-rater agreement between people, and low agreement between human raters and LLM-based methods. In response, we construct a personalization architecture for classifying value expressions by learning from a small number of highly informative calibration annotations per user. In evaluation, we find that modeling these differences successfully yields value expression predictions that people agree with more than they agree with other people. These results contribute new methods and understanding for the measurement of human values in social media data.
@@ Introduction
Social media provides a rich and complex substrate to understand patterns of human life, behavior, and discourse (Lazer et al. 2009). The rise of large language models (LLMs) offers new opportunities for measuring a wide variety of constructs in social media data (Ziems et al. 2024; Jia et al. 2024). One construct of interest for social media, growing out of the focus on value alignment of sociotechnical systems (Russell et al. 2015), is human values: the core beliefs that serve as guiding principles and in turn shape decisions, behavior, and perception (Kolluri et al. 2026; Bernstein et al. 2023; Stray et al. 2024). Values expressed on social media, therefore, offer a reflection of discourse and sentiment, that in turn affect behavior. The possibility of measuring human values in social media posts opens new opportunities to build a better understanding of how discourse operates online, and to design the next generation of social media algorithms and platforms. Past work has shown how values and related moral priorities explain patterns of belief and behavior in domains as varied as vaccine hesitancy (Weinzierl and Harabagiu 2022), pro-environmental behavior (Schultz and Zelezny 1998), social norms (Forbes et al. 2020), prosociality (Heilman and Kusev 2020), human development (Inglehart 2020), consumer behavior (Krystallis et al. 2012; Stathopoulou and Balabanis 2019), and news story framing (Mokhberian et al. 2020). Beyond understanding, the capability to measure values could also enable algorithmic objectives for feed-based ranking that support both individual and societal goals. In particular, there is growing interest in aligning ranking algorithms with societal values (Bail 2022; Bernstein et al. 2023; Stray et al. 2024; Kolluri et al. 2026), focusing on values such as democratic process (Jia et al. 2024; Piccardi et al. 2024a), well-being (Stray 2020), and downranking low-quality information (Epstein et al. 2020).
@@
Before these opportunities can be realized, there remain both conceptual and practical challenges in operationalizing and measuring value expressions in social media. By value expressions, we mean the extent to which a social media post expresses attributes associated with a given value. Identifying value expression can be straightforward in cases such as the post “Not all disabilities are visible!” (Figure 1), which many perceive as expressing the value of equality but not tradition.
@@
Figure 1: An anonymized example social media post for value annotation. All annotators in our dataset agree that this post strongly expresses Universal Concern, and they all agree that the post does not express Tradition—but, annotators disagree substantially on whether the post expresses Humility. In this paper, we model these interpersonal differences for more accurate value classification.
@@
However, not in all cases are value expressions as straightforward to classify on social media. The core issue is that value expression is a fundamentally subjective construct, leading to different labels from different people (Van Der Meer et al. 2023; Basile et al. 2021; Orlikowski et al. 2023). For example, while the Figure 1 post that “Not every disability is visible!” clearly is an expression of equality and clearly is not an expression of tradition, is it an expression of the value of humility? For some, the post strongly expresses a reminder to be humble about the conclusions we draw about other people’s (lack of) visible disabilities. For others, their perception of that value expression is much weaker. Both interpretations are valid and shaped by an individual’s background and how they both conceptualize these values and understand the social media post itself. As a result, we argue that there exists conceptual and methodological ambiguity about how value expression is operationalized. While existing work (e.g., Qiu et al. (2022); Van Der Meer et al. (2023); Borenstein et al. (2024)) has developed classifiers for detecting basic human values, these models have presumed there exists in all cases one unambiguous correct answer to be learned. But due to the possibility of divergent interpretations of value expressions, a traditional content-level ground truth set to train and evaluate models may obscure disagreements across individuals.
@@
In this paper, we therefore argue for a paradigm shift toward value labeling architectures that directly acknowledge the subjectivity of these constructs. In particular, we first identify disagreements in perceptions of values in social media posts, and then develop personalized methods to identify an individual’s perception of a value under a perspectivist framework. To do this, we collect a large annotated dataset where 1,079 people redundantly annotate values in 5,211 posts using a well-studied and theoretically grounded value system drawn from cross-cultural psychology — the Schwartz value system (Schwartz and Cieciuch 2022; Schwartz 1992). We evaluate to what extent large language models (LLMs) can identify the expression of values in these posts, comparing off-the-shelf LLM models with fine-tuned and personalized models, and both to individuals and groups of annotators.
@@
On this task, we find that not only are there low levels of agreement between people, but also that an off-the-shelf LLM exhibits ever lower levels of agreement with people than people’s levels of agreement with each other. To make sense of these findings, we draw on the theory of perspectivism — that different people will have different subjective experiences of the same situation (Soden et al. 2024; Berger and Luckmann 2016) — to suggest that people’s divergent perceptions of both the posts and values are driving the low levels of human-human agreement we observe, as well as the lack of performance of the off-the-shelf (base) GPT model.
With this idea in mind, we explicitly model interpersonal differences by combining two components: (1) fine-tuning a large language model to perform value inference, and (2) incorporating recommendation-system–inspired techniques that use a small, user-specific “cold-start” dataset to adapt predictions to individual users.
@@
We find that our personalized approach results in a performant model that individuals agree with more than the rate at which they agree with other people (66% relative improvement in spearman rank correlation ($\rho$)), or with the consensus vote (28% relative improvement in $\rho$).
We first contribute a demonstration that perceptions of value expression exhibit high levels of disagreement, which underscores the perils of relying on majority aggregated labels in this task. We then correspondingly contribute a theoretically-grounded personalization framework for calibrating predictions from a fine-tuned large language model to an individual based on their own responses. This results in predictions that align more closely with that person’s judgments than either other humans’ labels or few-shot–prompted LLM outputs.
@@
Figure 2: The refined Schwartz value system. The system is hierarchically nested into a highest level: Outcomes for Self (right) vs Outcomes for Others (left), a high level: Self-Transcendence (top left, green), Conservation (bottom left, blue), Self-Enhancement (bottom right, orange) and Openness to Change (top right, red), and the 19 low-level values. Adapted from (Schwartz et al. 2012)
@@
To operationalize values, we employ Schwartz’s theory of Basic Human Values (Schwartz et al. 2012), a theory from Cultural Psychology for mapping individual values. Schwartz defined (basic) values as “trans-situational goals, varying in importance, that serve as guiding principles in the life of a person or group” and sought to identify a set of values that could be recognized in all societies. The original theory (Schwartz 1992) identified 10 basic values from a 56-item questionnaire administered to 9140 students and teachers across 20 countries. The refined theory identified 19 values that exist on a circular continuum with 57-item questionnaire administered to 6,059 participants across 10 countries (see Figure 2). A core assumption of Schwartz’s theory of Basic Human Values is that: “The circular arrangement of values represents a continuum of related motivations, like the circular continuum of colors, rather than a set of discrete motivations” (Davidov et al. 2008). This circumplex structure is also reflected in the hierarchical branching structure of the values, which starts at the highest level (Outcomes for Self vs Others), then branches further to four high-level values (Openness to Change, Self-Enhancement, Self-Transcendence, and Conservation), and then 19 low-level values. Values close to each other on the wheel are theorized to have similar and compatible underlying motivations, while values on opposite sides are theorized to be in tension. While there are reasons to critique the Schwartz Value System (see e.g. (Chakroff 2015)), the theory ultimately offers conceptual precision, depth, and breadth, as well as offering more theoretical plausibility, analytical utility, and empirical grounding than political values (Goren 2020).
Detecting and Using Schwartz Values
@@
Qiu et al. (2022) create a dataset of human attitudes on over 20,000 text scenarios called ValueNet and a transformer-based regression model for dialogue tasks. Kiesel et al. (2022) create a dataset of the values behind arguments from four geographical cultures. Ponizovskiy et al. (2020) developed a dictionary of words for Schwartz values based on a corpus of Facebook updates, blog posts, essays, and book chapters. Boyd et al. (2015) use free-response survey methods to compare individual values reconstructed from text with traditional survey instruments, showing advancements in understanding in both the structure and content of values Van Der Meer et al. (2023) use data from Qiu et al. (2022) and Kiesel et al. (2022) to train a model for predicting value expression in Reddit posts, which they apply to a dataset of 11.4M comments from 19K users to construct value profiles of individual users. They find that dissimilarity of users’ values is associated with the levels of disagreement between those users. Borenstein et al. (2024) train value extraction models to classify both the presence and polarity of Schwartz values in reddit posts. They identify patterns like a strong negative stance towards conformity in subreddits such as r/Vegan and r/AbolishTheMonarchy an overall correlation between tradition values and the conservativeness of U.S. states. Shahid et al. (2025) use LLMs to re-write constructive comments on homophobic and Islamophobic threads, then uses human annotations to evaluate the values represented in these comments. They find that the comments rewritten by LLMs exhibited decreased Conservative values with increased prosocial values such as Benevolence and Universalism.
@@
This body of work uniformly makes the assumption that there exists an unambiguous “ground truth” value label, neglecting the inherent subjectivity of the task. In the next subsection, we discuss work that explores an alternative, perspectivist stance.
Perspectivism
@@
The work highlighted above on automatically detecting Schwartz values offers an exciting lens on the values represented in text. However, in every case, these models assume a fixed meaning for both the content and the values, and a corresponding unambiguous “objective” value label in each case. This is in line with the longstanding paradigm in data labeling (Fleisig et al. 2024) which seeks to collect “ground truth” labels (Nowak and Rüger 2010; Snow et al. 2008) by aggregating labels from annotators. An alternative, the perspectivist paradigm (Basile et al. 2021; Fleisig et al. 2024; Plank 2022; McDonald et al. 2019; Frenda et al. 2025) argues that variation in annotator labels is not purely a source of error to be expunged, but rather is intrinsically meaningful. In this vein, Haghighi et al. (2025) argues that the ontological dimensions underlying generative systems are under-recognized, and makes the case for perspectives that are plural (instead of universal), grounded (instead of abstract), lively (instead of fixed) and enacted (instead of diluted).
We argue that our task of identifying the expression of basic human values in social media posts fits this framework well, because the labels have no singularly correct “ground truth” label that everyone agrees with (they are plural) and are personally calibrated to each individual’s behavior (they are grounded).
Personalization
An alternative but complementary lens is personalization, which seeks to build models responsive to heterogeneous preferences. In the context of recommender systems, classic approaches involve matrix factorization (Koren et al. 2009) and collaborative filtering (Resnick et al. 1994). For social media algorithms, Seth and Zhang (2008) use a Bayesian user-model to learn personal recommender systems from social network data. Lerman (2006) uses social filtering of activities as a way to personalize recommendations. See Eg et al. (2023) for a review on social media personalization.
In the context of aligning LLMs, past work has explored aligning models to users’ opinions (Hwang et al. 2023; Do et al. 2025; Suh et al. 2025) and behaviors (Shaikh et al. 2025a). This is often task specific, with advances in domains as varied as computer usage (Shaikh et al. 2025b), creative writing (Chung et al. 2025) and user interface generation (Wu et al. 2025). In addition, recent work has explored personalized reward models (Chen et al. 2024; Jang et al. 2023; Li et al. 2024; Rame et al. 2023). This is achieved by conditioning of text-based user profiles (Zhang 2024; Ryan et al. 2025), using learned auxiliary predictors to steer towards group preferences (Zhao et al. 2023), and latent variable approaches (Poddar et al. 2024). In this work, we adopt a cold start approach (e.g. identifying and using a small yet compact set of past behaviors) because of its flexibility and interpretability.
@@
Methods
In this section, we first report our methods for curating a dataset of social media posts. Next, we discuss our methods for annotating this dataset for values. Finally, we provide details our approach to modeling and evaluating human and AI-based value expressions.
@@ Methods
We recruited a sample of social media users (N=281) quota-matched to be a representative sample of the US population on age, ethnicity, gender and partisanship using Prolific in May 2024. Participants were all from the United States and were required to be 18 years old or older. We were unable to meet the quota for 17/109 participants in the age range 55-100 so we upsample posts from participants in this age range by 18.4% when constructing our silverset of social media posts. Within this sample, the median age was 43 (min=19, max=80), 58.36% White/Caucasian, 51.2% female and 30.6% Democrat, 28.8% Republican and 40.6% Independent. Participants were paid $\sim\mathdollar 15.42$ an hour and took approximately 7 minutes to complete the study.
@@
After they consented to the research, participants were asked to install a Chrome Extension that downloaded posts from their X (Twitter) feeds, both their algorithmically curated For You page (FYP) and their in-network Following feeds simultaneously (Piccardi et al. 2024b). We collect and annotate posts from both the Following feed and FYP to account for any systematic differences that may exist between them, and to pave the way for work that uses both to measure algorithmic amplification (see Discussion).
When a browser opens the X page, tweets are fed to the browser client in “batches”: a single API request returns a set of tweets at a time. The Chrome Extension aims to download 16 batches of tweets: 7 from the FYP, and 9 from the Following page. Each batch for FYP consists of approximately 30 tweets, whereas the Following feed returns a variable number of tweets up to 100.
From the 281 participants, we collected 265,442 posts, with an average of 207 posts from the FYP and 776 from the Following page per user. Among these 265,442 posts, there are 151,740 posts with unique, non-empty text in English (using Google’s Chromium language detection tool cld2). We also screen out private posts (n=6,025; 3.97%), deleted posts (n=2,004; 1.32%) and/or reply tweets that are replying to posts not displayed in the user’s feed (n=7,581; 4.99%) to focus only on the 142,652 public posts we can render with proper context.
@@
We next prepare these posts for human annotation, since some posts are either not safe for work (NSFW) to show to annotators or incomprehensible to most annotators. We filter the posts for comprehensibility (grounded in the constructs of readability, coherence, spam behavior, and context required for understanding) and NSFW (to comply with our institution’s IRB for data annotation) using an automated GPT-based pipeline, (see Table A5 for full text and the Appendix for an evaluation of this model) with GPT-4o, temperature=1.0 and seed=0. We filter out any posts that score less than 3 on a comprehensibility scale, resulting in a dataset of 89,383 (62.6%) posts. We then filter out 1,937 posts with NSFW content, resulting in a final dataset of 87,446 posts.
@@
The occurrence rate of different values in social media posts varies, as do the numbers of posts per person across both Following and FYP feeds. Therefore, uniform sampling procedure may yield a sample that underrepresents rarer values and overrepresents certain users for annotation. To construct a final set of posts for human annotation, we stratify on user and post source (FYP vs following) as well as on a preliminary LLM-based screening measure of value expression, designed to err on the side of inclusion and capture as many plausibly value-expressing posts as possible.
@@
In particular, we perform a preliminary rough classification of these posts for all 19 of the Basic Human Values (Schwartz et al. 2012) on a 7-point Likert scale [0-6] using another automated GPT-based pipeline with the prompt that rates all 19 values jointly each on a 0-6 scale based on a short description, with the prompt defined in Table A6 with GPT-4o, temperature=1.0 and seed=0 (see Appendix for an evaluation of this method). This allows us to upsample posts that may reflect rarer values. For this upsampling, we say a given post reflects a given value if the GPT annotation pipeline scored it a 4 or above. We then sample a set of 5,227 posts for human annotation by stratifying across users, post source, and value expression. In particular, for each user and each value, we sample a tweet from that user’s feed that reflects that value. We sample FYP and Following posts with equal probability, but for a given user, if only one type of post exists, we return that tweet. Because we did not meet our quota for participants over the age of 55 (by 17 participants), we randomly sample (without replacement) 17 of the 76 participants over the age of 55 who completed the survey and repeat this process for each of them to ensure representation of posts from users over the age of 55. Of the 5,227, 3,182 came from users’ Following feeds (60.8%) and 2,045 (39.1%) came from users’ FYP. This process yielded a dataset of 5,227 posts that we can use to study human value annotation and to predict the values expressed in those posts.
Annotating Social Media Posts for Schwartz Values
@@
In this section, we describe our ground truth data annotation method.